Draw the molecular orbitals formed by the s – s combination of atomic orbitals. Also illustrate the
electron densities for these orbitals and explain them
1
Expert's answer
2020-04-15T01:50:59-0400

The drawing of molecular orbitals (MO) formed from s-s atomic orbitals (AO) is below. The axis shows increase of energy (E).

Electron densities are depicted as the clouds around the nuclei centered within them.

For the binding MO "\sigma" at lower energy the electron density homogeneously surrounds both atoms with the maximum in the interatomic region. For the antibonding MO "\sigma"* at higher energy electron density surrounds each atom, but has a minimum in the interatomic region.

Explanation: an orbital is a region with high enough probability of electron occurence (more than 90 %). The probability is a space integral of probability density, which is square of the absolute value of wavefunction "\psi" . The s-wavefunction is spherical and the atomic s-orbital is spherical too. The wavefunction corresponding to bonding MO is a linear sum of atomic s-wavefunctions: "\Psi(\sigma)=\psi1+\psi2" , that is why the probability (i.e. electron density) for bonding "\sigma" MO is proportional to the absolute value of "\Psi(\sigma)"2 = ("\psi1+\psi2")2 =

"\psi1"2 + 2*"\psi1"*"\psi2" + "\psi2"2. The sum ψ12 + ψ22 is a sum of electron densities of atomic s-orbitals, and the additive 2*ψ1*ψ2 is positive in the region where s-orbitals intercept, i.e. between atoms.

The wavefunction corresponding to antibonding MO is еру linear difference of atomic s-wavefunctions: Ψ(σ*)=ψ1 - ψ2 , that is why the electron density for antibonding σ* MO is proportional to the absolute value of Ψ(σ*)2 = (ψ1 - ψ2)2 = ψ12 - 2*ψ1*ψ2 + ψ22.

The sum ψ12 + ψ22 is a sum of electron densities of atomic s-orbitals, and the term (-2*ψ1*ψ2) is negative in the region where s-orbitals intercept, i.e. between atoms.

The green and violet colors show just the sign (+ or -) of atomic s-orbitals in the expressions for molecular orbitals.